# Chiral Heisenberg Gross-Neveu-Yukawa criticality: honeycomb vs. SLAC fermions

### v2

@@ Abstract
We perform large scale quantum Monte Carlo simulations of the Hubbard model at half filling with a single Dirac cone close to the critical point, which separates a Dirac semi-metal from an antiferromagnetically ordered phase where SU(2) spin rotational symmetry is spontaneously broken. We discuss the implementation of a single Dirac cone in the SLAC formulation for eight Dirac components and the influence of dynamically induced long-range super-exchange interactions. The finite size behavior of dimensionless ratios and the finite size scaling properties of the Hubbard model with a single Dirac cone are shown to be superior compared to the honeycomb lattice. We extract the critical exponent believed to belong to the chiral Heisenberg Gross-Neveu-Yukawa universality class: The critical exponent $\nu=1.02(3)$} coincides for the two lattice types once honeycomb lattices of linear dimension L>15{L>15} are considered. In contrast to the SLAC formulation, where the anomalous dimensions are estimated to be $\eta_{\phi}=0.73(1)$} and $\eta_{\psi}=0.09(1)$}, they remain less stable on honeycomb lattices, but tend towards the estimates from the SLAC formulation.
@@ I Introduction
The family of Gross-Neveu-Yukawa (GNY) universality classes capture the complex interplay of critical Dirac fermions coupled to a scalar bosonic fields. It represents the simplest universality of quantum critical points that lack classical analogues and realize the most fundamental scenario of relativistic fermions: the spontaneous acquisition of mass triggered by interactions [1, 2]. Here we focus specifically on the chiral Heisenberg [O(3) or SU(2)] transition [3] with N=8{N=8} Dirac components in (2+1)(2\!+\!1) dimensions found, e.g., in semi-metal to insulator transitions, which remains hard to control in analytical approaches [4, 5, 6, 7, 8, 9] as well as quantum Monte Carlo (QMC) simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Where the critical properties of the chiral Ising GNY transition have been arguably consistently determined to be at archipelago provided by conformal bootstrap [24], competing interaction channels and generically large corrections in finite size scaling prevent the same for the chiral Heisenberg fixed point [8].
The fairly disparate state of estimates for the N=8{N=8} critical exponents which allow us to classify the universality of the chiral Heisenberg transition is summarized in Fig. 1 along with the estimates from this investigation. In particular the results from QMC simulations, which rightfully pride themselves to be numerically exact, are far from a consensus. This illustrates the ambiguities in the finite size scaling analysis and the limitations by practically accessible spacetime volumes.
@@
Motivated by simulations in the Hamiltonian formulation of SLAC fermions [25, 26, 27, 28] in this manuscript we present the implementation and benchmark simulation results for the chiral-Heisenberg GNY quantum phase transition of the fermionic SLAC Hubbard model with $N=8$ Dirac components, which allows for a direct comparison with simulations on conventional lattices such as the honeycomb, or the $\pi$-flux lattice. We argue that the spread of the QMC results can be attributed to strong finite size corrections in conventional lattice implementations and show that our simulations, which maximize the linear Dirac dispersion across the Brillouin zone, significantly reduce finite size effects at the critical point. This allows for computationally favorable and more accurate estimates of the physics at criticality.
In the following Sec. II we describe the implementation and simulation details of congruent Dirac cones with a particular number $N$ of zero modes on the lattice and address the reliability and potential issues such as locality, renormalizability and dynamically induced non-local interactions in Sec. III. A finite size scaling analysis and estimates for the critical exponents are presented in Sec. IV. In the following we indicate results obtained with the SLAC formulation with a [Uncaptioned image] -icon, where the inscribed number denotes the number of massless modes $N$. Results obtained on the honeycomb lattice (N=8{N=8}) are identified with a [Uncaptioned image] -icon.
@@ II.1 Formulation for N-pole SLAC fermions
where the amplitudes t(r){t(r)} correspond to the Fourier amplitudes from the transformation of Eq. (1), which decay $\propto 1/r$} in the thermodynamic limit (TDL) $L\to\infty$}. Originally introduced for 3+1dd lattice field theories, SLAC fermions were constructed for odd $L=2N_{k}+1$} finite size systems [25], where
@@
SLAC fermions on lattices with even linear dimension $L=2N_{k}$}, have the amplitudes  [^1] Contrary to Ref. 52, lattices with even linear dimension do not induce Gibbs ringing, but also exhibit the perfect linear dispersion once the appropriate Fourier transformation for has been performed as shown in App. A.
@@
By construction, this perfectly reproduces the relativistic dispersion with a $2N_{f}$-fold degeneracy at $\mathbf{k}=(0,0)$} for $N_{f}$ fermion flavors, such that the eigenvalues at all finite size momenta obey ${\varepsilon_{\pm}(\mathbf{k})}=\pm v_{\text{F}}|\mathbf{k}|$, independently of the system size.
Figure 2: (a) Energy contours (light gray) and Berry flux pseudo spin map (blue arrows) for an $L=11$ lattice in momentum space. A closed contour around the Fermi point surface at the origin accumulates a flux of $\pi$ as can bee seen from the pseudo spin which sweeps twice the angle. The high symmetry path on a finite size system is indicated by the red dashed line. The dispersion shown in (b) is discontinuous at the boundary of the Brillouin zone.
@@ II.2 Method
A few technical drawbacks of the SLAC formulation compared to simulations on the honeycomb should be noted: (i) The numerical costs in AF-QMC simulations for the $N=8$ SLAC formulation with its four orbitals per unit cell scale as $O(64L^{6})$, which is higher than the $O(8L^{6})$ scaling of the honeycomb. As we will show though, the significantly smaller finite size effects make the SLAC formulation arguably still superior to the honeycomb implementation. This overhead could be overcome by a using a antisymmetic SU($N$) representation on the half-filled lattice, with an interaction term which solely breaks the targeted SU(2) symmetry, which we leave to future investigations. (ii) The non-local nature of the hopping term decreases the benefit employing a checkerboard breakup of the hopping matrix exponential. (iii) While we will show that the implementation with four fermion interactions works well, the combination of SLAC fermions with dynamical gauge fields constitutes not only a technical challenge, but also raises the questions with respect to the locality and renormalizability of such iteracting systems as described below. (iv) For even $L$ the last term in Eq. (4) corresponds to a constant amplitude, long range oscillating hopping, which will dynamically induce a homogeneous antiferromagnetic (AFM) magnetic field proportional to $\pi/L^{2}$} as discussed below. While it vanishes with system size, this symmetry breaking field will results in different finite size behavior compared to odd $L$ systems.
@@ III Locality & dynamically induced interactions
In this manuscript we simulate the Hamiltonian lattice model in Eq. (6), rather than the GNY lattice field theory itself. We have to be mindful of the apparent long range structure of the Dirac operator in real space, which stands in stark contrast to the inherent locality of physical quantum field theories, but allows to avoid the Nielsen-Ninomiya theorem [39, 40, 41] on finite size systems and to place a single Dirac cone onto the lattice  [^2] Note that in odd spacetime dimensions considered in this work chiral symmetry is replaced by parity symmetry [85].. The real-space hopping however is not equivalent to a genuine long range coupling, as the hopping only runs along the major axes. This couples a given site only to a finite fraction of the total volume. While this does not exclude possible spurious long-range effects, it suggests their potential influence scales intensively and decreases with increasing system size and dimensionality.
@@ III.2.1 Weak coupling Dirac semi-metal
We observe a stable DSM phase up to the critical point, beyond which AFM order emerges and the fermions acquire mass. The emergence of long-range AFM order can be tracked by measuring the spin structure factor $S_{\text{AFM}}(\mathbf{k})\equiv\sum_{\mathbf{r}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\langle\mathbf{S}(\mathbf{r})\cdot\mathbf{S}(\mathbf{0})\rangle/L^{2}$}, where the spin $\mathbf{S}_{\mathbf{r}}=\frac{1}{2}a_{\mathbf{r}\alpha}^{\dagger}\bm{\sigma}_{\alpha\beta}a_{\mathbf{r}\beta}$} at position $\mathbf{r}$, orbital $a$, and $\bm{\sigma}$ denotes the vector of the three Pauli matrices. The squared magnetization then is obtained from $m^{2}=S_{\text{AFM}}(\mathbf{k}=0)/L^{2}$}. Figure 3 shows the evolution of the finite size order parameter as a function of the Coulomb repulsion, which indicates the quantum phase transition to be located at $U_{c}\approx 6$}. The inset shows the finite size extrapolation of the single-particle gap $\Delta_{\text{sp}}(\mathbf{k})$ obtained from a fit to the asymptotic long imaginary-time behavior of the single particle Green’s function $G(\mathbf{k},\tau)=\langle a_{\mathbf{k}}^{\dagger}(\tau)\,a_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\text{sp}}(\mathbf{k})]$}. The gap vanishes faster than 1/L{1/L} consistent with simulations of local models on a torus [27, 59, 60] below the critical coupling $U\leq U_{c}$}, such that the Dirac fermions remain massless in the TDL and the gap opens with the onset of order. We detect no anisotropy in the single particle and spin Green’s functions, all supporting a stable semi-metal fixed point. Potential long-range spin fluctuations as a consequence of dynamically induced super-exchange mediated by the long range-hopping apparently do not destabilize the DSM.
@@
Figure 3: The model shows two well defined phases: a gapless DSM regime where the order parameter ($m^{2}$) vanishes and a gapped, AFM ordered phase where the order parameter acquires finite values. Inset: the finite size extrapolations of the single particle gap $\Delta_{\text{sp}}$ below, at and above the critical coupling $U_{c}\approx 6$}.
@@
The locality of the Dirac operator is recovered in the infinite volume limit for most of the BZ, but the singular boundary [61, 62, 44]. We thus monitor the dominant single particle channel and AFM correlations for interfering low-energy modes induced by interactions, in particular close to the boundary of the BZ. Such excitations (doublers) could potentially alter the degeneracy of the fermions at the Fermi level and consequently the universality class of the quantum phase transition. The single-particle spectral function $A(\mathbf{k},\omega)=-{\rm Im}\,G(\mathbf{k},\omega)/\pi$} close to the critical point as obtained from stochastic analytical continuation [63, 64] is presented in Fig. 4 for momenta along the path in the BZ sketched in Fig. 2(a). The dispersion of the non-interacting system is indicated by the white dashed line. The excitation gap at $\Gamma$ is the one most affected by finite size corrections, yet scales predominately as $\Delta_{\text{sp}}\propto L^{-z}$}. The finite size corrections of the single-particle excitations away from vanishing momentum are significantly smaller leading to a scaling faster than $L^{-1}$ [65, 60]. Note that the quasiparticle weight $Z$ (the intensity in Fig. 4) actually vanishes at $\mathbf{k}\to\Gamma$} at the critical point as $Z\sim(U-U_{c})^{-z\nu}$} in the TDL [27, 66]. The discontinuous dispersion at the BZ boundary appears to be smoothed upon the introduction of interactions and the single particle excitations close to the boundary of the BZ converge to a finite values at rather high energies. We indicate the finite size extrapolated values of the excitations at the momenta $M^{\prime}=(\pi-2\pi/L,0)$} and $X^{\prime}=(\pi-2\pi/L,\pi-2\pi/L)$} closest to the Brillouin zone boundary. No additional zero modes introduced by correlations as reflected in the interacting single-particle spectrum, nor the spin spectrum as discussed below at criticality.
@@
Figure 5: Estimators for the renormalization of the Fermi velocity based on the finite-size energies at the smallest lattice momentum $\Delta(\delta\mathbf{k})$, the relativistic envelope of $\Delta(\mathbf{k})$} and the Dirac point $\Delta(\mathbf{0})$, . The finite size show minor renormalization of the Fermi velocity.
@@
While the definition for $v_{\text{env}}$ includes the strongly finite size affected Dirac point, the use of SLAC fermions allows to reliably extract the excitations velocity from momenta where the dispersion is approximately linear. This is in contrast to the definition
@@ III.2.2 Strong coupling AFM
where we have used the explicit finite size expression for t(r){t(r)} from Eq. (3). The Hamiltonian (12) may be diagonalized by a Bogoliubov transformation, which yields the eigenvalues
@@
whereas for finite size systems we use $\varepsilon_{\text{LSWT}}(\mathbf{k})=\sqrt{J_{\mathbf{0}}^{2}-J_{\mathbf{k}}^{2}}$}. While retaining all lattice symmetries, the resulting dispersion is anisotropic as a consequence of the major axis coupling and shown in the inset of Fig. 6. In contrast, in fully connected long range models the magnon dispersion is isotropic. For small momenta the dispersion follows $\sqrt{k}$ similar to fully connected long range models [70], but for the major axis anisotropy. At strong coupling Eq. (16) approximates the true spin wave dispersion. This is shown for U=20{U=20}, N=4{N=4} along the cuts (k,0){(k,0)} and (k,k){(k,k)} in Fig. 7, where we extracted the spin gap $\Delta_{\sigma}$ obtained from a fit to the asymptotic long imaginary-time behavior of the unequal-time spin-spin correlation function $S(\mathbf{k},\tau)=\langle S_{\mathbf{k}}^{\dagger}(\tau)\,S_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\sigma}(\mathbf{k})]$}.
In the SU(2) spin symmetry broken phase one expects the emergence of Goldstone modes, which can be tested by the extrapolation $\Delta_{\sigma}(\mathbf{k}\to 0^{+})\to 0$}. Due to the nonlinear $\sqrt{k}$-form of the dispersion at small momenta, a naive extrapolation will significantly overestimate the gap $\Delta_{\sigma}(0^{+})$} in finite size systems with limited momentum resolution. In order to reduce the finite-size effects in the extrapolation, we rescale the finite-energy spectra at momentum $\mathbf{k}$ as
@@
where $v_{\phi}$ denotes the spin-wave velocity fitted to the data. This unwarping enables us to remove the leading non-linear behavior from the finite size data, yet retains the correct limit for vanishing momentum [60]. The then approximately linear dispersion allows for the extrapolation of the spin gaps to zero, which is consistent with the emergence of gapless Goldstone modes. While anomalous in their $\sqrt{k}$-dispersion, compared to a fully connected effective long-range Heisenberg model, the behavior also implies that the dynamically generated super-exchange coupling decays significantly faster than would be required for a gapped Goldstone mode [70] which has been claimed for the N=4{N=4} SLAC-Hubbard model in [53].
@@
Figure 7: The momentum dependence of the QMC spin gap $\Delta_{\sigma}(k)$ at strong coupling U=20{U=20} along the paths $(0,0)$ to $(\pi,0)$} and $(\pi,\pi)$} is well reproduced by the finite size LSWT (orange L=19{L=19}, gray $L\to\infty$}).
Figure 8: Finite size extrapolation the unwarped QMC data at strong coupling U=20{U=20} (c.f. inset) via the LSWT spin-wave dispersion $\varepsilon_{\text{LSWT}}$} to remove the leading nonlinear behavior. The linear finite size extrapolation of the smallest unwarped spin gaps $\hat{\Delta}_{\sigma}(k\to 0)$} to zero provides evidence for gapless, anomalous Goldstone modes.
The dynamically induced super-exchange along the major axis results in an anisotropic spin-wave dispersion accompanied by anomalous, yet gapless Goldstone modes in the AFM insulator. Despite the anisotropy, the AFM order still breaks the same SU(2) spin rotational symmetry while retaining the same lattice symmetries as in local interacting fermion models. We therefore argue that the we are considering the same strong coupling fixed point but for specific details of the spin correlations. Noticeably, we find no magnon excitations below the light cone at small momenta. This implies that the dynamical super-exchange does not induce instantaneous non-causal interactions, or only with vanishingly small amplitudes.

### v3

@@ Abstract
We perform large scale quantum Monte Carlo simulations of the Hubbard model at half filling with a single Dirac cone close to the critical point, which separates a Dirac semi-metal from an antiferromagnetically ordered phase where SU(2) spin rotational symmetry is spontaneously broken. We discuss the implementation of a single Dirac cone in the SLAC formulation for eight Dirac components and the influence of dynamically induced long-range super-exchange interactions. The finite-size behavior of dimensionless ratios and the finite-size scaling properties of the Hubbard model with a single Dirac cone are shown to be superior compared to the honeycomb lattice. We extract the critical exponent believed to belong to the chiral Heisenberg Gross-Neveu-Yukawa universality class: The critical exponent $\nu=1.02(3)$} coincides for the two lattice types once honeycomb lattices of linear dimension L>15{L>15} are considered. In contrast to the SLAC formulation, where the anomalous dimensions are estimated to be $\eta_{\phi}=0.73(1)$} and $\eta_{\psi}=0.09(1)$}, they remain less stable on honeycomb lattices, but tend towards the estimates from the SLAC formulation.
@@ I Introduction
The family of Gross-Neveu-Yukawa (GNY) universality classes capture the complex interplay of critical Dirac fermions coupled to a scalar bosonic fields. It represents the simplest universality of quantum critical points that lack classical analogues and realize the most fundamental scenario of relativistic fermions: the spontaneous acquisition of mass triggered by interactions [1, 2]. Here we focus specifically on the chiral Heisenberg [O(3) or SU(2)] transition [3] with N=8{N=8} Dirac components in (2+1)(2\!+\!1) dimensions found, e.g., in semi-metal to insulator transitions, which remains hard to control in analytical approaches [4, 5, 6, 7, 8, 9] as well as quantum Monte Carlo (QMC) simulations [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Where the critical properties of the chiral Ising GNY transition have been arguably consistently determined to be at archipelago provided by conformal bootstrap [24], competing interaction channels and generically large corrections in finite-size scaling prevent the same for the chiral Heisenberg fixed point [8].
The fairly disparate state of estimates for the N=8{N=8} critical exponents which allow us to classify the universality of the chiral Heisenberg transition is summarized in Fig. 1 along with the estimates from this investigation. In particular the results from QMC simulations, which rightfully pride themselves to be numerically exact, are far from a consensus. This illustrates the ambiguities in the finite-size scaling analysis and the limitations of practically accessible spacetime volumes.
@@
Motivated by simulations in the Hamiltonian formulation of SLAC fermions [25, 26, 27, 28] in this manuscript we present the implementation and benchmark simulation results for the chiral-Heisenberg GNY quantum phase transition of the fermionic SLAC Hubbard model with $N=8$ Dirac components, which allows for a direct comparison with simulations on conventional lattices such as the honeycomb, or the $\pi$-flux lattice. We argue that the spread of the QMC results can be attributed to strong finite-size corrections in conventional lattice implementations and show that our simulations, which maximize the linear Dirac dispersion across the Brillouin zone, significantly reduce finite-size effects at the critical point. This allows for computationally favorable and more accurate estimates of the physics at criticality.
In the following Sec. II we describe the implementation and simulation details of congruent Dirac cones with a particular number $N$ of zero modes on the lattice and address the reliability and potential issues such as locality, renormalizability and dynamically induced non-local interactions in Sec. III. A finite-size scaling analysis and estimates for the critical exponents are presented in Sec. IV. In the following we indicate results obtained with the SLAC formulation with a [Uncaptioned image] -icon, where the inscribed number denotes the number of massless modes $N$. Results obtained on the honeycomb lattice (N=8{N=8}) are identified with a [Uncaptioned image] -icon.
@@ II.1 Formulation for N-pole SLAC fermions
where the amplitudes t(r){t(r)} correspond to the Fourier amplitudes from the transformation of Eq. (1), which decay $\propto 1/r$} in the thermodynamic limit (TDL) $L\to\infty$}. Originally introduced for 3+1dd lattice field theories, SLAC fermions were constructed for odd $L=2N_{k}+1$} finite-size systems [25], where
@@
SLAC fermions on lattices with even linear dimension $L=2N_{k}$}, have the amplitudes  [^1] Contrary to Ref. 51, lattices with even linear dimension do not induce Gibbs ringing, but also exhibit the perfect linear dispersion once the appropriate Fourier transformation for has been performed as shown in App. A.
@@
By construction, this perfectly reproduces the relativistic dispersion with a $2N_{f}$-fold degeneracy at $\mathbf{k}=(0,0)$} for $N_{f}$ fermion flavors, such that the eigenvalues at all finite-size momenta obey ${\varepsilon_{\pm}(\mathbf{k})}=\pm v_{\text{F}}|\mathbf{k}|$, independently of the system size.
Figure 2: (a) Energy contours (light gray) and Berry flux pseudo spin map (blue arrows) for an $L=11$ lattice in momentum space. A closed contour around the Fermi point surface at the origin accumulates a flux of $\pi$ as can bee seen from the pseudo spin which sweeps twice the angle. The high symmetry path on a finite-size system is indicated by the red dashed line. The dispersion shown in (b) is discontinuous at the boundary of the Brillouin zone.
@@ II.2 Method
A few technical drawbacks of the SLAC formulation compared to simulations on the honeycomb should be noted: (i) The numerical costs in AF-QMC simulations for the $N=8$ SLAC formulation with its four orbitals per unit cell scale as $O(64L^{6})$, which is higher than the $O(8L^{6})$ scaling of the honeycomb. As we will show though, the significantly smaller finite-size effects make the SLAC formulation arguably still superior to the honeycomb implementation. This overhead could be overcome by using the antisymmetic SU($N$) representation on the half-filled lattice, with an interaction term which solely breaks the targeted SU(2) symmetry, which we leave to future investigations. (ii) The non-local nature of the hopping term decreases the benefit employing a checkerboard breakup of the hopping matrix exponential. (iii) While we will show that the implementation with four fermion interactions works well, the combination of SLAC fermions with dynamical gauge fields constitutes not only a technical challenge, but also raises the questions with respect to the locality and renormalizability of such iteracting systems as described below. (iv) For even $L$ the last term in Eq. (4) corresponds to a constant amplitude, long range oscillating hopping, which will dynamically induce a homogeneous antiferromagnetic (AFM) magnetic field proportional to $\pi/L^{2}$} as discussed below. While it vanishes with system size, this symmetry breaking field will results in different finite-size behavior compared to odd $L$ systems.
@@ III Locality & dynamically induced interactions
In this manuscript we simulate the Hamiltonian lattice model in Eq. (6), rather than the GNY lattice field theory itself. We have to be mindful of the apparent long range structure of the Dirac operator in real space, which stands in stark contrast to the inherent locality of physical quantum field theories, but allows to avoid the Nielsen-Ninomiya theorem [38, 39, 40] on finite-size systems and to place a single Dirac cone onto the lattice  [^2] Note that in odd spacetime dimensions considered in this work chiral symmetry is replaced by parity symmetry [84].. The real-space hopping however is not equivalent to a genuine long range coupling, as the hopping only runs along the major axes. This couples a given site only to a finite fraction of the total volume. While this does not exclude possible spurious long-range effects, it suggests their potential influence scales intensively and decreases with increasing system size and dimensionality.
@@ III.2.1 Weak coupling Dirac semi-metal
We observe a stable DSM phase up to the critical point, beyond which AFM order emerges and the fermions acquire mass. The emergence of long-range AFM order can be tracked by measuring the spin structure factor $S_{\text{AFM}}(\mathbf{k})\equiv\sum_{\mathbf{r}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\langle\mathbf{S}(\mathbf{r})\cdot\mathbf{S}(\mathbf{0})\rangle/L^{2}$}, where the spin $\mathbf{S}_{\mathbf{r}}=\frac{1}{2}a_{\mathbf{r}\alpha}^{\dagger}\bm{\sigma}_{\alpha\beta}a_{\mathbf{r}\beta}$} at position $\mathbf{r}$, orbital $a$, and $\bm{\sigma}$ denotes the vector of the three Pauli matrices. The squared magnetization then is obtained from $m^{2}=S_{\text{AFM}}(\mathbf{k}=0)/L^{2}$}. Figure 3 shows the evolution of the finite-size order parameter as a function of the Coulomb repulsion, which indicates the quantum phase transition to be located at $U_{c}\approx 6$}. The inset shows the finite-size extrapolation of the single-particle gap $\Delta_{\text{sp}}(\mathbf{k})$ obtained from a fit to the asymptotic long imaginary-time behavior of the single particle Green’s function $G(\mathbf{k},\tau)=\langle a_{\mathbf{k}}^{\dagger}(\tau)\,a_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\text{sp}}(\mathbf{k})]$}. The gap vanishes faster than 1/L{1/L} consistent with simulations of local models on a torus [27, 58, 59] below the critical coupling $U\leq U_{c}$}, such that the Dirac fermions remain massless in the TDL and the gap opens with the onset of order. We detect no anisotropy in the single particle and spin Green’s functions, all supporting a stable semi-metal fixed point. Potential long-range spin fluctuations as a consequence of dynamically induced super-exchange mediated by the long range-hopping apparently do not destabilize the DSM.
@@
Figure 3: The model shows two well defined phases: a gapless DSM regime where the order parameter ($m^{2}$) vanishes and a gapped, AFM ordered phase where the order parameter acquires finite values. Inset: the finite-size extrapolations of the single particle gap $\Delta_{\text{sp}}$ below, at and above the critical coupling $U_{c}\approx 6$}.
@@
The locality of the Dirac operator is recovered in the infinite volume limit for most of the BZ, but the singular boundary [60, 61, 43]. We thus monitor the dominant single particle channel and AFM correlations for interfering low-energy modes induced by interactions, in particular close to the boundary of the BZ. Such excitations (doublers) could potentially alter the degeneracy of the fermions at the Fermi level and consequently the universality class of the quantum phase transition. The single-particle spectral function $A(\mathbf{k},\omega)=-{\rm Im}\,G(\mathbf{k},\omega)/\pi$} close to the critical point as obtained from stochastic analytical continuation [62, 63] is presented in Fig. 4 for momenta along the path in the BZ sketched in Fig. 2(a). The dispersion of the non-interacting system is indicated by the white dashed line. The excitation gap at $\Gamma$ is the one most affected by finite-size corrections, yet scales predominately as $\Delta_{\text{sp}}\propto L^{-z}$}. The finite-size corrections of the single-particle excitations away from vanishing momentum are significantly smaller leading to a scaling faster than $L^{-1}$ [64, 59]. Note that the quasiparticle weight $Z$ (the intensity in Fig. 4) actually vanishes at $\mathbf{k}\to\Gamma$} at the critical point as $Z\sim(U-U_{c})^{-z\nu}$} in the TDL [27, 65]. The discontinuous dispersion at the BZ boundary appears to be smoothed upon the introduction of interactions and the single particle excitations close to the boundary of the BZ converge to a finite values at rather high energies. We indicate the finite-size extrapolated values of the excitations at the momenta $M^{\prime}=(\pi-2\pi/L,0)$} and $X^{\prime}=(\pi-2\pi/L,\pi-2\pi/L)$} closest to the Brillouin zone boundary. No additional zero modes introduced by correlations as reflected in the interacting single-particle spectrum, nor the spin spectrum as discussed below at criticality.
@@
Figure 5: Estimators for the renormalization of the Fermi velocity based on the finite-size energies at the smallest lattice momentum $\Delta(\delta\mathbf{k})$, the relativistic envelope of $\Delta(\mathbf{k})$} and the Dirac point $\Delta(\mathbf{0})$, . The finite-size show minor renormalization of the Fermi velocity.
@@
While the definition for $v_{\text{env}}$ includes the strongly finite-size affected Dirac point, the use of SLAC fermions allows to reliably extract the excitations velocity from momenta where the dispersion is approximately linear. This is in contrast to the definition
@@ III.2.2 Strong coupling AFM
where we have used the explicit finite-size expression for t(r){t(r)} from Eq. (3). The Hamiltonian (12) may be diagonalized by a Bogoliubov transformation, which yields the eigenvalues
@@
whereas for finite-size systems we use $\varepsilon_{\text{LSWT}}(\mathbf{k})=\sqrt{J_{\mathbf{0}}^{2}-J_{\mathbf{k}}^{2}}$}. While retaining all lattice symmetries, the resulting dispersion is anisotropic as a consequence of the major axis coupling and shown in the inset of Fig. 6. In contrast, in fully connected long range models the magnon dispersion is isotropic. For small momenta the dispersion follows $\sqrt{k}$ similar to fully connected long range models [69], but for the major axis anisotropy. At strong coupling Eq. (16) approximates the true spin wave dispersion. This is shown for U=20{U=20}, N=4{N=4} along the cuts (k,0){(k,0)} and (k,k){(k,k)} in Fig. 7, where we extracted the spin gap $\Delta_{\sigma}$ obtained from a fit to the asymptotic long imaginary-time behavior of the unequal-time spin-spin correlation function $S(\mathbf{k},\tau)=\langle S_{\mathbf{k}}^{\dagger}(\tau)\,S_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\sigma}(\mathbf{k})]$}.
In the SU(2) spin symmetry broken phase one expects the emergence of Goldstone modes, which can be tested by the extrapolation $\Delta_{\sigma}(\mathbf{k}\to 0^{+})\to 0$}. Due to the nonlinear $\sqrt{k}$-form of the dispersion at small momenta, a naive extrapolation will significantly overestimate the gap $\Delta_{\sigma}(0^{+})$} in finite-size systems with limited momentum resolution. In order to reduce the finite-size effects in the extrapolation, we rescale the finite-energy spectra at momentum $\mathbf{k}$ as
@@
where $v_{\phi}$ denotes the spin-wave velocity fitted to the data. This unwarping enables us to remove the leading non-linear behavior from the finite-size data, yet retains the correct limit for vanishing momentum [59]. The then approximately linear dispersion allows for the extrapolation of the spin gaps to zero, which is consistent with the emergence of gapless Goldstone modes. While anomalous in their $\sqrt{k}$-dispersion, compared to a fully connected effective long-range Heisenberg model, the behavior also implies that the dynamically generated super-exchange coupling decays significantly faster than would be required for a gapped Goldstone mode [69] which has been claimed for the N=4{N=4} SLAC-Hubbard model in [52].
@@
Figure 7: The momentum dependence of the QMC spin gap $\Delta_{\sigma}(k)$ at strong coupling U=20{U=20} along the paths $(0,0)$ to $(\pi,0)$} and $(\pi,\pi)$} is well reproduced by the finite-size LSWT (orange L=19{L=19}, gray $L\to\infty$}).
Figure 8: finite-size extrapolation the unwarped QMC data at strong coupling U=20{U=20} (c.f. inset) via the LSWT spin-wave dispersion $\varepsilon_{\text{LSWT}}$} to remove the leading nonlinear behavior. The linear finite-size extrapolation of the smallest unwarped spin gaps $\hat{\Delta}_{\sigma}(k\to 0)$} to zero provides evidence for gapless, anomalous Goldstone modes.
The dynamically induced super-exchange along the major axis results in an anisotropic spin-wave dispersion accompanied by anomalous, yet gapless Goldstone modes in the AFM insulator. Despite the anisotropy, the AFM order still breaks the same SU(2) spin rotational symmetry while retaining the same lattice symmetries as in local interacting fermion models. We therefore argue that we are considering the same strong coupling fixed point but for specific details of the spin correlations. Noticeably, we find no magnon excitations below the light cone at small momenta. This implies that the dynamical super-exchange does not induce instantaneous non-causal interactions, or only with vanishingly small amplitudes.